Determine the area under the standard normal distribution curve between z=1.15 and z=-1.85

1
Expert's answer
2022-01-30T15:29:22-0500

P(-1.85<z<1.15)

P(z<-1.85) = 0.032157

P(z<1.15) = 0.8749

P(z<1.15)- p(z<-1.85) = 0.8749-0.0322

P(z<1.15)- p(z<-1.85)=0.84277